Two concentric conducting loops have radii Ra and Rb (Ra>>Rb). The inner loop has
current I flowing counter-clockwise through it, and the outer loop has no current. The outer
loop has resistance R. The inner loop is cut by scissors, so that in a fairly small duration
current in it drops to zero. What is the direction of induced current in the outer loop? How
much total charge passes a point in the loop during the period this current flows?
1
Expert's answer
2020-04-28T09:46:39-0400

According to Lenz's law, the current in the outer loop will try to generate a magnetic field that would prevent the field in the inner loop to cease. Therefore, according to the figure below and the right-hand rule, the current in the outer loop will flow counter-clockwise as well.


According to Faraday's and Ohm's law:


"\Epsilon=\frac{\Delta \Phi}{\Delta t}=iR,\\
\space\\
i\Delta t=q,\\
\space\\
\frac{\Delta \Phi}{\Delta t}=\frac{q}{\Delta t}R.\\
\space\\
q=\frac{\Delta \Phi}{R}."

The inner loop creates the flux that disappears of


"\Delta\Phi=BA=\frac{\mu_0I}{2R_a}\pi R_a^2=\frac{\pi\mu_0IR_a}{2}."

Substitute this into the expression for the charge:


"q=\frac{\pi\mu_0IR_a}{2R}."